# Inland-LOAM: Voxel-Based Structural Semantic Mapping for Inland Waterways

### v1

@@ 4.2 Evaluation of Trajectory and Point Cloud Mapping
To highlight the performance of our method, we conducted comparative experiments against several state-of-the-art methods: KISS-ICP, HDL-SLAM, LeGO-LOAM, and LOAM. To ensure a fair and consistent comparison, all methods were adapted to match the laser beam distribution specific to our dataset.
@@
To assess the quality of the generated point cloud maps, we compared Inland-LOAM visually with LOAM, LeGO-LOAM, HDL-SLAM, and KISS-ICP within Area B, as illustrated in Fig. 10. The gradient coloring along the trajectory indicates elevation changes. Notably lacking bridges or large overhead structures, Area B presents a challenge for evaluating vertical localization accuracy.
@@
At a global scale, LOAM and KISS-ICP show a water surface that curves downwards along the z-axis, while HDL-SLAM exhibits an upward curvature. In contrast, LeGO-LOAM displays only minor z-axis drift. Inland-LOAM maintained a relatively flat water surface, demonstrating superior mitigation of vertical drift. At a local scale, Fig. 10 details the mapping quality of the same building produced by each method, reflecting the precision of their horizontal registration. KISS-ICP, LOAM, and LeGO-LOAM fail to reconstruct the building’s windows, capturing only the main facade. While HDL-SLAM correctly identifies the windows, it fails to represent their proper alignment. Inland-LOAM, however, accurately reconstructs both the windows and their correct spatial arrangement, a clear benefit of our optimized feature extraction algorithm.
@@
The accuracy of the trajectory estimation was quantitatively assessed by comparing Inland-LOAM with LOAM, LeGO-LOAM, KISS-ICP, and HDL-SLAM. Fig. 11 presents these comparative trajectories. We calculated the Root Mean Square Error (RMSE) and Standard Deviation (STD) between RTK-GNSS ground truth positions and the trajectories generated by these methods, as summarized in Table 2. It is important to note that, although RTK-GNSS data were treated as ground truth, GNSS signals can experience drift in certain inland waterway scenarios, particularly under bridges or in narrow urban sections in Figs. 11(a) (sections 1 and 3) and (c) (right side).
@@
As shown in Fig. 11(b), the Inland-LOAM trajectory closely overlaps the ground truth of RTK-GNSS, while the other methods exhibit a noticeable gap. In Figs. 11(a) and  11(c), HDL-SLAM displayed a larger gap, while Inland-LOAM consistently maintained a minimal deviation from the ground truth. As shown in Table 2, Inland-LOAM achieves the lowest RMSE and standard deviation in each test case. Overall, the trajectory results indicate that although all methods performed similarly in general localization within inland waterways, Inland-LOAM achieved superior accuracy and robustness under challenging conditions. This high-precision trajectory estimation provides a solid foundation for subsequent semantic analysis and mapping tasks.
@@
Figure 11: Comparison of the trajectories of each algorithm in experiments. Sub-figures (a-2), (b), and the left of (c) provide magnified views for examination of local accuracy. Sub-figures (a-1), (a-3), and the right of (c) project the trajectories from Inland-LOAM (blue) and the ground truth (black) onto IENC to demonstrate accuracy in confined spaces.
Table 2: Algorithm trajectory accuracy evaluation.
Inland-LOAM HDL-SLAM KISS-ICP LeGO-LOAM LOAM
Seq RMSE STD RMSE STD RMSE STD RMSE STD RMSE STD
A 0.45 0.37 3.14 1.34 1.28 0.64 0.48 0.37 1.27 0.58
B 0.60 0.30 4.87 2.11 9.61 4.78 2.26 1.14 7.44 3.59
C 0.99 0.72 2.25 1.06 2.22 1.08 2.01 0.83 1.97 0.97
Avr 0.68 0.46 3.42 1.50 4.37 2.17 1.58 0.78 3.56 1.71
4.3 Structural Semantic Mapping Evaluation
@@
Figure 14: Structural semantic maps generated by Inland-LOAM in areas A, B, and C. Each scenario is presented with corresponding real-world photographs from multiple perspectives, annotated to explain key features. The color encoding corresponds to the semantic classes defined in Section 3.4.2: red for Horizontal Planar Surfaces (HP), pink for Vertical Planar Surfaces (VP), green for Complex Horizontal Terrain (CH), yellow for Complex Vertical Structures (CV), blue for Open Water Surface (OW), and black for Under Bridge Clearance Zone (UB). Note that this color scheme is used for visualization purposes only and does not represent the traditional cost definitions in a costmap. For (a) and (c), detailed semantic maps of bridge areas are included, with the identified clearance height explicitly labeled (H).
@@
Fig. 14 presents the 2D structural semantic maps generated for our three experimental areas A, B, and C. Fig. 12(a) shows the performance along an inland waterway with highly diverse shoreline characteristics. The system accurately identifies navigable areas under bridges, providing real-time bridge clearance estimations (H-1, H-2, H-3), which are essential for the safe passage of large vessels and address the limitations of traditional static IENCs. Furthermore, it demonstrates the ability to differentiate shoreline constructions, distinguishing boat ramps classified as HP at c-1, shore platforms (HP) at e-3, containers classified as VP at d-2, and bridge piers (VP) at e-1. It also distinguishes vegetated banks classified as CH and CV at a-2, and piling identified as CV at c-2.
Fig. 13(a) illustrates a scenario dominated by dense reed vegetation. The map correctly captures the overall characteristics of the canal, categorizing large buildings at location b-1 as VP and most of the shoreline as complex terrain CH or CV, aligning well with real-world scenes. However, this scenario highlights limitations of purely geometric classification. Scattered small HP and VP regions appear within extensively classified vegetation areas (e.g., near c-2 and d-2). Such ambiguity likely results from LiDAR penetrating sparse foliage and capturing flat ground or mistaking dense, flat-topped vegetation as planar surfaces. Although this reveals the inherent challenges of using geometric methods in complex natural environments, it also demonstrates the ability to detect subtle variations in surface structure.
In Fig. 14(a), the system successfully identifies bridges and labels the navigable water beneath UB. It clearly classifies large building facades in b-3 as VP, and precisely outlines vessels (b-2, c-2) and revetments (b-1, c-1). However, movement of the lifting bridge (H-1) during data acquisition led to color mixing between the UB and VP areas, revealing limitations of our system in dynamic environments.
@@
4.4 Shoreline Boundaries Comparison with Reference Maps
@@
In Area A (Fig. 15(a)), shoreline boundaries (red hits) align with the contours of artificial structures such as revetments and shoreline constructions, exhibiting strong correspondence with the IENC and Google Earth. In contrast, for vegetated banks, some discrepancies appear relative to the satellite imagery, reflecting seasonal variation in reeds or minor erosion not represented in the static maps. In Area B (Fig. 15(b)), our system accurately delineates boundaries, including vessels moored long-term along the canal. Area C (Fig. 15(c)) demonstrates a distinct case. Although the canal boundary aligns with all reference datasets, the layout of moored vessels within the port basin shows a notable divergence. The positions identified by Inland-LOAM are consistent with the latest updates in OSM and differ from the outdated layouts present in other map, reflecting Inland-LOAM’s potential to dynamically enhance and maintain up-to-date navigational charts.
@@
4.5 Time analysis
Table 3 evaluates the real-time performance of the LOAM module. The front-end components, which include pre-processing, plane fitting, feature extraction, and LiDAR odometry, have a combined processing time of 18.3 ms. The back-end mapping module is 37.2 ms. Both timings are shorter than the data acquisition interval required by the LiDAR’s sampling frequency (10 Hz) and the map generation frequency (2 Hz), respectively, confirming the real-time performance of our LOAM module.
For the conversion module, the computational time required to transform the voxel map into the 2D structural semantic map depicted in Fig. 16. The voxel map update time scales with map size, and this time cost is particularly pronounced in open environments compared to enclosed scenes [40]. In our experiments, Area A is the only trajectory that includes loop closures, resulting in significantly lower computation times compared to areas B and C, which are focused on open area exploration. This difference arises because in loop-back scenarios, only the updated map segments require modification, while open-area exploration triggers broader map updates. Despite this, the median map conversion time for areas B and C is approximately 1s, with a maximum of 2.8s. These update times are sufficient within the requirements for real-time decision making and control on a USV driving at up to 3m/s, given a 150m LiDAR sensing range.
The IENC integration module, primarily designed for long-term maintenance and sharing of known environment maps, is executed offline. This approach ensures that the generated IENC-compatible data maintains high accuracy and stability without imposing additional computational cost during navigation tasks.
Table 3: Processing Time for the Components of LOAM Module.
@@ 4.5 Time analysis
5 Conclusion
In this paper, we proposed Inland-LOAM, an integrated LiDAR-only SLAM framework specifically designed for robust odometry, mapping, and structural semantic map generation in challenging inland waterway environments. Addressing the limitations of conventional SLAM methods, our contributions include an enhanced feature extraction approach and a joint optimization method that uses the water surface as a globally consistent planar reference to effectively reduce drift. Furthermore, we detailed an end-to-end pipeline transforming dense 3D point clouds into structured, semantic 2D maps. The proposed voxel-based semantic segmentation and subsequent 2D mapping method allow real-time estimation of critical navigational parameters, such as bridge clearances, and provide precisely delineated shorelines compatible with IENC standards.
@@ 5 Conclusion
6 CRediT authorship contribution statement
Zhongbi Luo: Conceptualization, Methodology, Investigation, Resources, Validation, Visualization, Formal analysis, Software, Data curation, Writing – original draft, Writing – review & editing. Yunjia Wang: Writing – review & editing, Data curation. Jan Swevers: Writing – review & editing. Peter Slaets: Project administration, Supervision, Funding acquisition, Writing – review & editing. Herman Bruyninckx: Conceptualization, Methodology, Visualization, Supervision, Writing – review & editing.
7 Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
8 Acknowledgment
We would like to thank the subject editor and anonymous reviewers who provided insightful and constructive comments on the manuscript. This project has received funding from the European Union’s EU Framework Programme for Research and Innovation Horizon 2020 under Grant Agreement No. 955768 (ETN AUTOBarge). Project website: https://etn-autobarge.eu/.

### v2

@@ IV-B Evaluation of Trajectory and Point Cloud Mapping
To highlight the performance of our method, we conducted comparative experiments against several state-of-the-art methods: KISS-ICP, HDL-SLAM, LeGO-LOAM, and LOAM. To ensure a fair and consistent comparison, all methods were adapted to match the laser beam distribution specific to our dataset. Specifically for HDL-SLAM, we configured it for odometry-only with planar constraints enabled.
@@
To assess the quality of the generated point cloud maps, we compared Inland-LOAM visually with LOAM, LeGO-LOAM, HDL-SLAM, and KISS-ICP within Area B, as illustrated in Figure 9. The gradient coloring along the trajectory indicates elevation changes. Notably lacking bridges or large overhead structures, Area B presents a challenge for evaluating vertical localization accuracy.
@@
At a global scale, LOAM and KISS-ICP show a water surface that curves downwards along the z-axis, while HDL-SLAM exhibits an upward curvature. In contrast, LeGO-LOAM displays only minor z-axis drift. Inland-LOAM maintains a relatively flat water surface, demonstrating superior mitigation of vertical drift. At a local scale, Figure 9 details the mapping quality of the same building produced by each method, reflecting the precision of their horizontal registration. KISS-ICP, LOAM, and LeGO-LOAM fail to reconstruct the building’s windows, capturing only the main facade. While HDL-SLAM correctly identifies the windows, it fails to represent their proper alignment. However, Inland-LOAM accurately reconstructs both the windows and their correct spatial arrangement, a clear benefit of our optimized feature extraction algorithm.
@@
The accuracy of the trajectory estimation was quantitatively assessed by comparing Inland-LOAM with LOAM, LeGO-LOAM, KISS-ICP, and HDL-SLAM. Figure 10 presents these comparative trajectories. We calculated the Root Mean Square Error (RMSE) and Standard Deviation (STD) between RTK-GNSS ground truth positions and the trajectories generated by these methods, as summarized in Table II. It is important to note that, although RTK-GNSS data were treated as ground truth, GNSS signals can experience drift in certain inland waterway scenarios, particularly under bridges or in narrow urban sections in Figure 10(a) (sections 1 and 2) and Figure 10(c) (right side).
@@
As shown in Figure 10(b), the Inland-LOAM trajectory closely aligns with the RTK-GNSS ground truth, while other methods exhibit a noticeable gap. In Figure 10(a) and  10(c), HDL-SLAM and KISS-ICP display a larger gap, while Inland-LOAM consistently maintains a minimal deviation from the ground truth. Regarding the Z-axis comparison, Inland-LOAM demonstrates the lowest error in all scenarios. Notably, in Figure 10(a), the Z-axis trajectory of LeGO-LOAM is similar to that of Inland-LOAM, unlike significant errors from other scenarios. The primary reason is that Area A contains multiple bridges; those overhead structures provide strong geometric observations that reduce the degeneracy of the Z-axis estimation. As shown in Table II, Inland-LOAM achieves the lowest RMSE and standard deviation in each test case. In particular, in Area A, the loop closure module had a negative effect, resulting in a higher RMSE compared to the odometry-only version. We attribute this to the highly similar scenery of the narrow waterway, which likely caused false positive loop closures, introducing errors that outweighed the accumulated odometry drift. Overall, the trajectory results indicate that although all methods performed similarly in general localization within inland waterways, Inland-LOAM achieved superior accuracy and robustness under challenging conditions. This high-precision trajectory estimation provides a solid foundation for subsequent semantic analysis and mapping tasks.
@@
Figure 10: Trajectory comparison of all evaluated algorithms against the ground truth. The main plots compare the estimated trajectories against the ground truth in both horizontal and vertical dimensions, with each trajectory only aligned with the ground truth at the origin point. Sub-figures (a-1), (a-2), and the right of (c) project the Inland-LOAM (blue) and ground truth (black) trajectories onto IENC to demonstrate accuracy in confined spaces, after rigidly aligning Inland-LOAM to the ground truth.
TABLE II: Algorithm trajectory accuracy evaluation.
Inland-LOAM Inland-LOAM* HDL-SLAM HDL-SLAM* KISS-ICP LeGO-LOAM LeGO-LOAM* LOAM
Seq RMSE STD RMSE STD RMSE STD RMSE STD RMSE STD RMSE STD RMSE STD RMSE STD
A 0.45 0.37 0.49 0.39 3.13 1.34 3.14 1.34 1.28 0.64 0.48 0.37 0.77 0.55 1.27 0.58
B 0.60 0.30 - - 4.87 2.11 - - 9.61 4.78 2.26 1.14 - - 7.44 3.59
C 0.99 0.72 - - 2.25 1.06 - - 2.22 1.08 2.01 0.83 - - 1.97 0.97
Avr 0.68 0.46 - - 3.42 1.50 - - 4.37 2.17 1.58 0.78 - - 3.56 1.71
*Variant with the loop closure module activated.
IV-C Structural Semantic Mapping Evaluation
@@
Figure 13: Structural semantic maps generated by Inland-LOAM in areas A, B, and C. Each scenario is presented with corresponding real-world photographs from multiple perspectives, annotated to explain key features. The color encoding corresponds to the semantic classes defined in Section III-D2: red for Horizontal Planar Surfaces (HP), pink for Vertical Planar Surfaces (VP), green for Complex Horizontal Terrain (CH), yellow for Complex Vertical Structures (CV), blue for Open Water Surface (OW), and black for Under Bridge Clearance Zone (UB). Note that this color scheme is used for visualization purposes only and does not represent the traditional cost definitions in a costmap. For (a) and (c), detailed semantic maps of bridge areas are included, with the identified clearance height explicitly labeled (H).
@@
Figure 13 presents the 2D structural semantic maps generated for our three experimental areas A, B, and C. Figure 11(a) shows the performance along an inland waterway with highly diverse shoreline characteristics. The system accurately identifies navigable areas under bridges, providing real-time bridge clearance estimations (H-1, H-2, H-3), which are essential for the safe passage of large vessels and address the limitations of traditional static IENCs. Furthermore, it demonstrates the ability to differentiate shoreline constructions, distinguishing boat ramps classified as HP at c-1, shore platforms (HP) at e-3, containers classified as VP at d-2, and bridge piers (VP) at e-1. It also distinguishes vegetated banks classified as CH and CV at a-2, and piling identified as CV at c-2.
Figure 12(a) illustrates a scenario dominated by dense reed vegetation. The map correctly captures the overall characteristics of the canal, categorizing large buildings at location b-1 as VP and most of the shoreline as complex terrain CH or CV, aligning well with real-world scenes. However, this scenario highlights limitations of purely geometric classification. Scattered small HP and VP regions appear within extensively classified vegetation areas (e.g., near c-2 and d-2). Such ambiguity likely results from LiDAR penetrating sparse foliage and capturing flat ground or mistaking dense, flat-topped vegetation as planar surfaces. Although this reveals the inherent challenges of using geometric methods in complex natural environments, it also demonstrates the ability to detect subtle variations in surface structure.
In Figure 13(a), the system successfully identifies bridges and labels the navigable water beneath UB. It clearly classifies large building facades in b-3 as VP, and precisely outlines vessels (b-2, c-2) and revetments (b-1, c-1). However, movement of the lifting bridge (H-1) during data acquisition led to color mixing between the UB and VP areas, revealing limitations of our system in dynamic environments.
@@
IV-D Shoreline Boundaries Comparison with Reference Maps
@@
In Area A (Figure 14(a)), shoreline boundaries (red hits) align with the contours of artificial structures such as revetments and shoreline constructions, exhibiting strong correspondence with the IENC and Google Earth. In contrast, for vegetated banks, some discrepancies appear relative to the satellite imagery, reflecting seasonal variation in reeds or minor erosion not represented in the static maps. In Area B (Figure 14(b)), our system accurately delineates boundaries, including vessels moored long-term along the canal. Area C (Figure 14(c)) demonstrates a distinct case. Although the canal boundary aligns with all reference datasets, the layout of moored vessels within the port basin shows a notable divergence. The positions identified by Inland-LOAM are consistent with the latest updates in OSM and differ from the outdated layouts present in other map, reflecting Inland-LOAM’s potential to dynamically enhance and maintain up-to-date navigational charts.
@@
IV-E Time analysis
Table III evaluates the real-time performance of the LOAM module. The front-end components, which include pre-processing, plane fitting, feature extraction, and LiDAR odometry, have a combined processing time of 18.3 ms. The back-end mapping module is 37.2 ms. Both timings are shorter than the data acquisition interval required by the LiDAR’s sampling frequency (10 Hz) and the map generation frequency (2 Hz), respectively, confirming the real-time performance of our LOAM module.
For the conversion module, the computational time required to transform the voxel map into the 2D structural semantic map depicted in Figure 15. The voxel map update time scales with map size, and this time cost is particularly pronounced in open environments compared to enclosed scenes [41]. In our experiments, Area A is the only trajectory that includes loop closures, resulting in significantly lower computation times compared to areas B and C, which are focused on open area exploration. This difference arises because in loop-back scenarios, only the updated map segments require modification, while open-area exploration triggers broader map updates. Despite this, the median map conversion time for areas B and C is approximately 1s, with a maximum of 2.8s. These update times are sufficient within the requirements for real-time decision making and control on a USV driving at up to 3m/s, given a 150m LiDAR sensing range.
The IENC integration module, primarily designed for long-term maintenance and sharing of known environment maps, is executed offline. This approach ensures that the generated IENC-compatible data maintain high accuracy and stability without imposing additional computational cost during navigation tasks.
TABLE III: Processing Time for the Components of LOAM Module.
@@ IV-E Time analysis
V Conclusion
In this paper, we propose Inland-LOAM, an integrated LiDAR-only SLAM framework specifically designed for robust odometry, mapping, and structural semantic map generation in challenging inland waterway environments. Addressing the limitations of conventional SLAM methods, our contributions include an enhanced feature extraction approach and a joint optimization method that uses the water surface as a globally consistent planar reference to effectively reduce drift. Furthermore, we detailed an end-to-end pipeline transforming dense 3D point clouds into structured, semantic 2D maps. The proposed voxel-based semantic segmentation and subsequent 2D mapping method allow real-time estimation of critical navigational parameters, such as bridge clearances, and provide precisely delineated shorelines compatible with IENC standards.
@@ V Conclusion
Acknowledgment
The authors would like to thank the Editor-in-Chief, an Associate Editor, and anonymous referees for their invaluable comments and suggestions.